# Diffusion Random Feature Model

### v2

@@ 2.1 Forward Process
The forward process degrades the input data such that $q({\mathbf{x}}_{K})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$. More precisely, let $\mathbf{x}_{0}\in{\mathbb{R}}^{d}$ be an input from an unknown distribution with p.d.f. $q({\mathbf{x}}_{0})$. Given a variance schedule $10<\beta_{1}<\beta_{2}<...<\beta_{K}<$ 1, the forward process to obtain a degraded sample for a given timestep is defined as:
@@
[…]
@@ 2.2 Reverse Process
$=\dfrac{\sqrt{(1-\overline{\alpha}_{k})^{d}}}{\sqrt{(2\pi\beta_{k}(1-\overline{\alpha}_{k-1}))^{d}}}\exp\left(-\frac{1}{2}\dfrac{({\mathbf{x}}_{k-1}-\tilde{\boldsymbol{\mu}}_{k})^{T}({\mathbf{x}}_{k-1}-\tilde{\boldsymbol{\mu}}_{k})}{\tilde{\beta}_{k}}\right)$ (5)
@@ 3 Algorithm and Theory
1:Sample ${\mathbf{x}}_{0}\sim q({\mathbf{x}}_{0})$ where $q$ is an unknown distribution, variance schedule $\beta=\{\beta_{1},...,\beta_{K}\}$ such that $10<\beta_{1}<\beta_{2}<...<\beta_{K}<$ 1, random weight matrix $\mathbf{W}=[w_{ij}]$ and bias vector $\mathbf{b}$ sampled from a distribution $\rho$ and total number of epochs epoch.
@@
7:Define the forward process for $k=1,2,...,K$ as
@@
20:     $\tilde{{\mathbf{x}}}_{k-1}=\dfrac{1}{\sqrt{1-\beta_{k}}}\left({\mathbf{x}}_{k}-\dfrac{\sqrt{\beta_{k}}}{\sqrt{1-\prod\limits_{i=1}^{t}(1-\beta_{i})}}p_{\theta}({\mathbf{x}}_{k},k))\right)+\beta_{k}{\boldsymbol{\epsilon}}$
@@ 3.1 Theoretical Results
For each $j=1,...,N$, let $a_{j}=\sin\left(\displaystyle\sum\limits_{i=1}^{d}y_{i}\omega_{ij}+b_{j}\right)$. Note that $a_{j}$’s are fixed. Then Eq. (21) becomes,
@@
[…]
If $[\boldsymbol{\omega}_{j}]_{j\in[N]}$ are drawn iid from $\rho$, then for $\delta>0$, with probability at least $1-\delta$ over $[\boldsymbol{\omega}_{j}]_{j\in[N]}$, there exists a function $\boldsymbol{f}^{\sharp}\in\mathcal{F}_{\boldsymbol{\omega}}$ such that
@@
[…]
@@ 4 Experimental Results
We compare our method with: (1) fully connected version of DRFM denoted by NN where we train $[\mathbf{W},\mathbf{b},\boldsymbol{\theta}^{(1)},\boldsymbol{\theta}^{(2)}]$ while preserving the number of trainable parameters; (2) classical random feature approach with only $\boldsymbol{\theta}^{(2)}$ being trainable denoted by RF. The details of the implementation of all the experiments and their results are described in the sections below.

### v3

# Generalization Bound for Diffusion Models using Random Features

@@ 2.1 Forward Process
The forward process degrades the input data such that $q({\mathbf{x}}_{K})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$. More precisely, let $\mathbf{x}_{0}\in{\mathbb{R}}^{d}$ be an input from an unknown distribution with p.d.f. $q({\mathbf{x}}_{0})$. Given a variance schedule $0<\beta_{1}<\beta_{2}<...<\beta_{K}<1$, the forward process to obtain a degraded sample for a given timestep is defined as:
@@
where ${\boldsymbol{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Since $0<\beta_{1}<...<\beta_{K}<1$, $0<\overline{\alpha}_{K}<\alpha_{1}^{K}<1$. Therefore, $\lim\limits_{K\rightarrow\infty}\overline{\alpha}_{K}=0$. Hence, $q({\mathbf{x}}_{K})=\int q({\mathbf{x}}_{K}|{\mathbf{x}}_{0})q({\mathbf{x}}_{0})d{\mathbf{x}}_{0}\approx\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e., as the number of timesteps becomes very large, the distribution $q({\mathbf{x}}_{K})$ will approach the Gaussian distribution with mean $\mathbf{0}$ and covariance $\mathbf{I}$.
@@ 2.2 Reverse Process
$\dfrac{\sqrt{(1-\overline{\alpha}_{k})^{d}}}{\sqrt{(2\pi\beta_{k}(1-\overline{\alpha}_{k-1}))^{d}}}\exp\left(-\frac{1}{2}\dfrac{({\mathbf{x}}_{k-1}-\tilde{\boldsymbol{\mu}}_{k})^{T}({\mathbf{x}}_{k-1}-\tilde{\boldsymbol{\mu}}_{k})}{\tilde{\beta}_{k}}\right)$ (5)
@@ 3 Algorithm and Theory
1:Sample ${\mathbf{x}}_{0}\sim q({\mathbf{x}}_{0})$ where $q$ is an unknown distribution, variance schedule $\beta=\{\beta_{1},...,\beta_{K}\}$ such that $0<\beta_{1}<\beta_{2}<...<\beta_{K}<1$, random weight matrix $\mathbf{W}=[w_{ij}]$ and bias vector $\mathbf{b}$ sampled from a distribution $\rho$ and total number of epochs epoch.
@@
7:Define the forward process for k=1,2,,Kk=1,2,...,K as
@@
20:  $\tilde{{\mathbf{x}}}_{k-1}=\dfrac{1}{\sqrt{1-\beta_{k}}}\left({\mathbf{x}}_{k}-\dfrac{\sqrt{\beta_{k}}}{\sqrt{1-\prod\limits_{i=1}^{k}(1-\beta_{i})}}p_{\theta}({\mathbf{x}}_{k},k))\right)+\beta_{k}{\boldsymbol{\epsilon}}$
@@ 3.1 Theoretical Results
For each j=1,,Nj=1,...,N, let $a_{j}=\sin\left(\displaystyle\sum\limits_{i=1}^{d}y_{i}\omega_{ij}+b_{j}\right)$. Note that $a_{j}$’s are fixed. Then Eq. (21) becomes,
@@
$\mathcal{F}_{\rho}=\left\{\boldsymbol{f}(\mathbf{x})=\int_{\Omega}\boldsymbol{\alpha}(\boldsymbol{\omega})\phi(\mathbf{x};\boldsymbol{\omega})\,d\boldsymbol{\omega}\ \Bigg{|}\ \|\boldsymbol{\alpha}(\boldsymbol{\omega})\|_{\infty}\leq C{\rho(\boldsymbol{\omega})}\right\}.$
If $\boldsymbol{\omega}_{j}]_{j\in[N]}$ are drawn iid from $\rho$, then for $\delta>0$, with probability at least $1-\delta$ over $\boldsymbol{\omega}_{j}]_{j\in[N]}$, there exists a function $\boldsymbol{f}^{\sharp}\in\mathcal{F}_{\boldsymbol{\omega}}$ such that
@@
We follow the proof technique described in [23]. As ${\boldsymbol{f}}^{*}\in\mathcal{F}_{\rho}$, then ${\boldsymbol{f}}^{*}({\mathbf{x}})=\displaystyle\int_{\boldsymbol{\Omega}}\boldsymbol{\alpha}(\boldsymbol{\omega})\phi({\mathbf{x}};\boldsymbol{\omega})d\boldsymbol{\omega}$. Construct ${\boldsymbol{f}}_{k}=\boldsymbol{\beta}_{k}\phi(\cdot;\boldsymbol{\omega}_{k})$, $k=1,\cdots,N$ such that $\boldsymbol{\beta}_{k}=\dfrac{\boldsymbol{\alpha}(\boldsymbol{\omega}_{k})}{\rho(\boldsymbol{\omega}_{k})}=\dfrac{1}{\rho(\boldsymbol{\omega}_{k})}\begin{bmatrix}\alpha_{1}(\boldsymbol{\omega}_{k})\\ \vdots\\ \alpha_{d}(\boldsymbol{\omega}_{k})\end{bmatrix}$.
Note that $\mathbb{E}_{\boldsymbol{\omega}}({\boldsymbol{f}}_{k})=\displaystyle\int_{\boldsymbol{\omega}}\dfrac{\boldsymbol{\alpha}(\boldsymbol{\omega}_{k})}{\rho(\boldsymbol{\omega}_{k})}\phi({\mathbf{x}};\boldsymbol{\omega})\rho(\boldsymbol{\omega}_{k})d\boldsymbol{\omega}={\boldsymbol{f}}^{*}$.
@@ 4 Experimental Results
We compare our method with: (1) fully connected version of DRFM denoted by NN where we train $\mathbf{W},\mathbf{b},\boldsymbol{\theta}^{(1)},\boldsymbol{\theta}^{(2)}]$ while preserving the number of trainable parameters; (2) classical random feature approach with only $\boldsymbol{\theta}^{(2)}$ being trainable denoted by RF. The details of the implementation of all the experiments and their results are described in the sections below.